OpenStudy (coolguy34):

step by step solution please give me the step by step solution to this question : 2x + 5 – 4x = -3

pooja195 (pooja195):

(2x+−4x)+(5)=−3(Combine Like Terms) −2x+5=−3 −2x+5=−3

pooja195 (pooja195):

Subtract 5 from both sides. −2x+5−5=−3−5 −2x=−8

pooja195 (pooja195):

Divide both sides by -2

pooja195 (pooja195):

-8/-2=?

OpenStudy (gabebae):

Nevermind it looks like you got it

OpenStudy (coolguy34):

thanks @pooja195

pooja195 (pooja195):

np